# Post-Newtonian templates for phase evolution of spherical extreme mass ratio inspirals

### v2

@@ Abstract
We present various post-Newtonian (PN) models for the phase evolution of compact objects moving along quasi-spherical orbits in Kerr spacetime derived by using the 12PN analytic formulas of the energy, angular momentum and their averaged rates of change calculated in the framework of the black hole perturbation theory. To examine the convergence of time-domain PN models (TaylorT families), we evaluate the dephasing between approximants with different PN orders. We found that the TaylorT1 model shows the best performance and the performance of the TaylorT2 is the next best. To evaluate the convergence of frequency-domain PN models (TaylorF families), we evaluate the mismatch between approximants with different orders. We found that the performance of the TaylorF2 model is comparable with the TaylorT2 model. Although the TaylorT2 and TaylorF2 models are not so accurate as the TaylorT1, the fully analytical expressions give us easy-to-handle templates and are useful to discuss effects beyond general relativity.
@@ I Introduction
Matched filtering is one of promising techniques to analyze the observed GWs and extract the physical information from signals emitted by EMRIs accurately and efficiently. This technique requires waveform templates representing the real signals from EMRIs accurately. The GW templates of EMRIs are often calculated by black hole perturbation theory because the extreme mass ratio is suitable for the small parameter in the perturbative expansion.
@@
In the test-mass limit, the motion is expressed by a geodesic in Kerr spacetime. Taking into account the self-interaction with the own gravitational field, however, a deviation of the mass’ motion from the geodesic arises. This can be interpreted as the “self-force” (see Refs. [10, 11] and the references therein) effect. The dominant contribution of the gravitational self-force to the mass’ motion is the time-averaged dissipative piece at the first order of the mass ratio, which induces the losses of the energy and angular momentum of the point mass and drives the gradual inspiral into the central black hole. The dissipative effect is most noticeable on the orbital phase and therefore the GW phase. A phase difference of $O(1)$ significantly reduces the overlap between an observed signal and the corresponding template, and leads to failure of the detection or the accurate estimation of the physical parameters. Therefore, it is important to calculate the time-averaged dissipative self-force at high precision.
@@
The time-averaged dissipative self-force can be rewritten in the secular changes of three orbital parameters, the energy, azimuthal angular momentum, and Carter parameter of the point mass. There are several works on calculating the rates of the secular change in both numerical and analytical approaches based on the black hole perturbation theory. The state-of-the-art studies on the numerical approach can provide the data of the secular evolution for general bound orbits in Kerr spacetime [12, 13, 14, 15]. Since the numerical calculation costs a lot of computational resources, however, the use of combination with the analytical approach is expected to reduce the cost and improve the efficiency. There have been several efforts in analytically calculating the secular evolution of the orbital parameters since the 1990s (see Ref. [16, 17] and the references therein). The post-Newtonian (PN) formulas of the secular evolution for general bound orbits are presented in our previous works [18, 19, 20]. The recent advances in computers have made it possible to extend the analytical calculation to the higher PN order. Several works on calculating the higher PN corrections have been done in Refs. [21, 22, 23, 24, 25].
@@
In this work, we focus on the secular evolution of orbits off equatorial plane in Kerr spacetime, which has attracted less attention in the study of the analytic calculation of the self-force so far. As the first step, we consider the evolution of a spherical orbit, which is an orbit with a constant radius on a plane precessing around the rotation axis of the central black hole. We calculate the 12PN formulas for the secular change of orbital parameters for spherical orbits with arbitrary inclination angle, through the procedure based on the adiabatic approximation proposed in Refs. [19, 20]  [^1] In Refs. [19, 20], the PN and eccentricity-expanded formulas for generic bound orbits not only spherical orbits are presented. Since the higher PN calculation with finite eccentricity requires significantly more cost than the non-eccentric orbit case, however, we restrict the calculation to the spherical orbit case in this work..
@@
To study the convergence of the PN formulas and the impact of the higher corrections on the accuracy of the gravitational waveform, we derive the PN templates of the phase evolution, the time-domain, TaylorT and frequency-domain, TaylorF models [27, 28], for quasi-spherical orbits. Although the TaylorT and TaylorF templates in the context of the black hole perturbation theory have been derived in Ref. [29], the work is restricted to the circular orbit case in Schwarzschild spacetime in exchange for the remarkably high, 22PN order. This work presents the extension into the spherical orbit case in Kerr spacetime, and we investigate how the convergence of the PN formulas is influenced by the orbital inclination and the spin of the central black hole.
@@
This paper is organized as follows. In Sec. II, we briefly review the spherical orbit in Kerr spacetime and its secular evolution driven by the dissipation due to the gravitational radiation. In Sec. III, we present the TaylorT and TaylorF models derived in a similar manner to Ref. [29]. In addition, we propose alternative choices of the TaylorT1 and TaylorT4 templates, which appear because of increasing the number of orbital parameters. In Sec. IV, we examine the performance of the Taylor models for several sets of the orbital inclination and the black hole spin. In Sec. V, we summarize our results. In Appendices AI, we present the explicit expressions of the PN coefficients of the orbital energy and angular momentum, their time averaged rates of change, the TaylorT and TaylorF templates, up to the 4.5PN order. Throughout this paper, we adopt the geometrized units with $c=G=1$ and metric signature ($+$)(-$+$)( - + + + ), and the Kerr metric is described in the Boyer–Lindquist coordinates $(t,r,\theta,\varphi)$.
@@ II.1 Geodesic equations in Kerr spacetime
Here, $z=\cos\theta$. The geodesic motion is characterized by three parameters, $I^{A}=\{E,L,C\}$, corresponding to the specific energy, angular momentum and Carter constant, respectively. For bound orbits, we can choose an alternative parametrization analogous to celestial mechanics, $\iota^{A}=\{p,e,\chi\}$, defined by
$p\equiv\frac{2r_{a}r_{p}}{M(r_{a}+r_{p})},\quad e\equiv\frac{r_{a}-r_{p}}{r_{a}+r_{p}},\quad\cos\chi\equiv\frac{L}{\sqrt{L^{2}+C}},$ (9)
where $r_{p}$ and $r_{a}$ are the values of $r$ at the periapsis and apoapsis. We refer $p$, $e$, and $\chi$ to the non-dimensional semi-latus rectum, eccentricity, and inclination angle, respectively.
@@
The adiabatic evolution of $\iota^{A}$ and orbital phase (i.e., rotational angle) $\Phi$ are described by
$\dot{\iota}^{A}$ $=$ $(G^{-1})^{A}_{B}\dot{I}^{B},$ (15)
@@
where a dot ($\cdot$) over variable denotes derivative with respect to $t$ and $G^{A}_{B}\equiv(\partial I^{A}/\partial\iota^{B})$ is the Jacobian for the transformation from $\iota^{A}$ to $I^{A}$.
@@ II.2 Spherical orbits
A spherical orbit is an bound orbit with $e=0$ (or equivalently, $r_{p}=r_{a}$), and hence is characterized by two orbital parameters, $\{p,\chi\}$. The parameters, $\{E,L,C\}$, are not independent for spherical orbits. In this work, we choose $\{E,L\}$ from them as a set of independent parameters.
For the later convenience, we introduce $x\equiv(M\omega_{\varphi})^{1/3}$ and $Y\equiv\cos\chi$ instead of $p$ and $\chi$ respectively. Since $x$ corresponds to the characteristic velocity in the motion, it is used as the parameter of the PN expansion. By using the set of $\{x,Y\}$, the equations of orbital evolution, Eqs. (15) and (16), for spherical orbits, are reexpressed as
@@
The specific energy and angular momentum for a spherical orbit at the leading order of the mass ratio are given in the form of the PN expansion as
@@
The superscript $(n)$ in the left hand side of the above equations means that the formula includes the terms up to $n$-th order terms of $x$ relative to the Newtonian order  [^2] We consider the Newtonian order of the energy is $E^{(0)}=1-x^{2}/2$. The term like $x^{n}(\ln x)^{k}$ is treated as $O(x^{n})$.. For example, the 2.5PN ($n=5$) energy and angular momentum are
$E^{(5)}(x,Y)$ $=$ $1-\frac{1}{2}x^{2}\left\{1-\frac{3}{4}x^{2}-\left(\frac{4}{3}-4Y\right)qx^{3}-\left[\frac{27}{8}-\left(\frac{1}{2}+Y-\frac{5}{2}Y^{2}\right)q^{2}\right]x^{4}+(2+6Y)qx^{5}\right\},$
$L^{(5)}(x,Y)$ $=$ $\frac{Y}{x}\left\{1+\frac{3}{2}x^{2}+\left(\frac{2}{3}-4Y\right)x^{3}+\left[\frac{27}{8}-\left(\frac{1}{4}+\frac{Y}{2}-\frac{7}{4}Y^{2}\right)q^{2}\right]x^{4}-(1+6Y)x^{5}\right\},$
@@
where $\mu$ denotes the mass of a point mass. In the same way of Eqs. (20) and (21), the superscript $(n)$ means the truncation order. It should be noted that the coefficients, $\tilde{\dot{E}}_{k}$ and $\tilde{\dot{L}}_{k}$, include some terms with the power of $\ln x$ for $k\geq 6$. In Appendix B, we show $\tilde{\dot{E}}_{k}$ and $\tilde{\dot{L}}_{k}$ for $k\leq 9$ explicitly.
@@ III Post-Newtonian templates of the phase evolution
The PN approximation provides various expressions of the phase evolution with different variables and functional forms while keeping the same PN order [27, 28, 29]. In this section, we present several PN families of the phase evolution of quasi-spherical orbits, TaylorT1, TaylorT2, TaylorT3, and TaylorT4 in the time domain  [^3] We do not consider the TaylorEt approximant because it is shown that the performance is not good compared to the other PN families for non-spinning, circular case [28, 29]. In Kerr case, the specific energy depends on the odd power of $x$ and therefore the expansion of $x$ with respect to $\zeta=-2E$ used in TaylorEt includes the half-integer powers, which makes the approximant more complicated. This is another reason for not considering the TaylorEt in this paper., and TaylorF1, TaylorF2 in the frequency domain, in a similar manner to Refs. [27, 28, 29]. In addition, we present subfamilies of TaylorT1 and TaylorT4, which appear because the number of the orbital parameters increases compared to circular orbits.
@@ III.1 TaylorT1
Substituting the PN formulas given by Eqs. (20)–(23) to $E$, $L$, $\dot{E}$ and $\dot{L}$ in the right hand sides of Eqs. (17) and (18), we obtain
@@
$\dot{\Phi}_{\textrm{T1}}^{(n)}$ $=$ $\frac{x^{3}}{M},$ (26)
where $(G_{n})^{A}_{B}=\partial(E^{(n)},L^{(n)})/\partial(x,Y)$. Note that we do not perform any expansion with respect to $x$ after the substitution. Therefore, the right hand sides of Eqs. (24) and (25) can be expressed by rational functions of $x$, $\ln x$ and $Y$ although we do not show them explicitly because of their complexity.
By analogy of the circular orbit case, we refer the numerical solution of Eqs. (24)–(26) as the TaylorT1 approximant.
@@ III.2 TaylorT4
Expanding the right hand sides of Eqs. (24) and (25) with respect to $x$, we obtain
@@
It should be noted that, since the PN formulas of the first and second terms in the right hand side of Eq. (25) are given as
@@
the first two leading terms of them are canceled each other out. Hence, the leading term in the PN expansion of $\dot{Y}$ is considered as the 1.5PN contribution of the adiabatic evolution.
The TaylorT4 approximant is obtained by solving Eqs. (27)–(31) numerically.

### v3

@@ Abstract
We present various post-Newtonian (PN) models for the phase evolution of compact objects moving along quasi-spherical orbits in Kerr spacetime. The models are derived by using the 12PN analytic formulas of the energy, angular momentum and their averaged rates of change calculated in the framework of the black hole perturbation theory. To examine the convergence of time-domain PN models (TaylorT families), we evaluate the dephasing between approximants with different PN orders. We found that the TaylorT1 model shows the best performance and the performance of the TaylorT2 model is the next best. To evaluate the convergence of frequency-domain PN models (TaylorF families), we evaluate the mismatch between approximants with different orders. We found that the performance of the TaylorF2 model is comparable with the TaylorT2 model. Although the TaylorT2 and TaylorF2 models are not so accurate as the TaylorT1 model, the fully analytical expressions give us easy-to-handle templates and are useful to discuss effects beyond general relativity.
@@ I Introduction
Matched filtering is one of the promising techniques to analyze the observed GWs and extract the physical information from signals emitted by EMRIs accurately and efficiently. This technique requires waveform templates representing the real signals from EMRIs accurately. The GW templates of EMRIs are often calculated by black hole perturbation theory because the extreme mass ratio is suitable for the small parameter in the perturbative expansion.
@@
In the test-mass limit, the motion is expressed by a geodesic in Kerr spacetime. Taking into account the self-interaction with the own gravitational field, however, a deviation of the mass’ motion from the geodesic arises. This can be interpreted as the “self-force” (see Refs. [10, 11] and the references therein) effect. The dominant contribution of the gravitational self-force to the mass’ motion is the time-averaged dissipative piece at the first order of the mass ratio. It induces the losses of energy and angular momentum of the point mass and drives the gradual inspiral into the central black hole. The dissipative effect is most noticeable on the orbital phase and therefore the GW phase. A phase difference of $O(1)$ significantly reduces the overlap between an observed signal and the corresponding template, and leads to failure of the detection or the accurate estimation of the physical parameters. Therefore, it is important to calculate the time-averaged dissipative self-force at high precision.
@@
The time-averaged dissipative self-force can be rewritten in the secular changes of three orbital parameters, the energy, azimuthal angular momentum, and Carter parameter of the point mass. There are several works on calculating the rates of the secular change in both numerical and analytical approaches based on the black hole perturbation theory. The state-of-the-art studies on the numerical approach can provide the data of the secular evolution for general bound orbits in Kerr spacetime [12, 13, 14, 15]. Since the numerical calculation costs a lot of computational resources, however, the use of combination with the analytical approach is expected to reduce the cost and improve the efficiency. There have been several efforts in analytically calculating the secular evolution of the orbital parameters since the 1990s (see Refs. [16, 17] and the references therein). The post-Newtonian (PN) formulas of the secular evolution for general bound orbits are presented in our previous works [18, 19, 20].
The recent advances in computers have made it possible to extend the analytical calculation to the higher PN order. Several works on calculating the high PN (8th or higher PN) corrections have been done in Refs. [21, 22, 23, 24, 25]. We can also find works on the high PN formulas for the self-force correction to the redshift invariant [26, 27, 28, 29]. The high PN calculations so far, however, are limited to equatorial orbit cases. On the other hand, the PN calculations for non-equatorial orbits remain at the lower order than those for equatorial cases. We find the 2.5PN results for circular orbits with small inclination [30] and for eccentric orbits with small inclination [18] (see also Ref. [19] with arbitrary inclination angle), the 4PN results for generic bound orbits [20], which have been recently updated to 5PN order [31]. According to the major scenarios of EMRI formation [32, 33], many EMRIs are expected to be not only eccentric but also inclined. Therefore, it is natural for one to be interested in extending the PN calculation for non-equatorial orbits to the higher order  [^1] In fact, at the same time as we submitted this paper, another work on the calculation of the 12PN fluxes for spherical orbits [34] had appeared independently..
@@
In this work, as a first step of the high PN calculation for non-equatorial orbits, we focus on the secular evolution of a spherical orbit, which is an orbit with a constant radius on a plane precessing around the rotation axis of the central black hole. Recently, a calculation of self-forced inspirals for spherical orbits has been done in numerical approach [35]. In contrast, we here calculate the 12PN formulas for the secular change of orbital parameters through the analytic procedure based on the adiabatic approximation proposed in Refs. [19, 20]  [^2] In Refs. [19, 20], the PN and eccentricity-expanded formulas for generic bound orbits not only spherical orbits are presented. Since the higher PN calculation with finite eccentricity requires significantly more cost than the non-eccentric orbit case, however, we restrict the calculation to the spherical orbit case in this work..
@@
To study the convergence of the PN formulas and the impact of the higher corrections on the accuracy of the gravitational waveform, we derive the PN templates of the phase evolution, the time-domain, TaylorT and frequency-domain, TaylorF models [36, 37], for quasi-spherical orbits. Although the PN templates in the context of the black hole perturbation theory have been derived in Ref. [38], the work is restricted to the circular orbit case in Schwarzschild spacetime in exchange for the remarkably high, 22PN order. This work presents the extension into the spherical orbit case in Kerr spacetime, and we investigate how the convergence of the PN formulas is influenced by the orbital inclination and the spin of the central black hole.
@@
This paper is organized as follows. In Sec. II, we briefly review the spherical orbit in Kerr spacetime, and its secular evolution driven by the dissipation due to the gravitational radiation. In Sec. III, we present the TaylorT and TaylorF models derived in a similar manner to Ref. [38]. In addition, we propose alternative choices of the TaylorT1 and TaylorT4 templates, which appear because of increasing the number of orbital parameters. In Sec. IV, we examine the performance of the Taylor models for several sets of the orbital inclination and the black hole spin. In Sec. V, we summarize our results. In Appendices AI, we present the explicit expressions of the PN coefficients of the orbital energy and angular momentum, their time averaged rates of change, the TaylorT and TaylorF templates, up to the 4.5PN order (the higher order terms are publicly available online [31]). Throughout this paper, we adopt the geometrized units with $c=G=1$ and metric signature ($+$)(-$+$)( - + + + ), and the Kerr metric is described in the Boyer–Lindquist coordinates $(t,r,\theta,\varphi)$.
@@ II.1 Geodesic equations in Kerr spacetime
Here, $z=\cos\theta$. The geodesic motion is characterized by three parameters, $I^{A}=\{E,L,C\}$, corresponding to the specific energy, angular momentum and Carter constant, respectively. For bound orbits, we can choose an alternative parametrization analogous to celestial mechanics, $\chi^{A}=\{p,e,\iota\}$, defined by
$p\equiv\frac{2r_{a}r_{p}}{M(r_{a}+r_{p})},\quad e\equiv\frac{r_{a}-r_{p}}{r_{a}+r_{p}},\quad\cos\iota\equiv\frac{L}{\sqrt{L^{2}+C}},$ (9)
where $r_{p}$ and $r_{a}$ are the values of $r$ at the periapsis and apoapsis. We refer $p$, $e$, and $\iota$ to the non-dimensional semi-latus rectum, eccentricity, and inclination angle, respectively.
@@
The adiabatic evolution of $\chi^{A}$ and orbital phase (i.e., rotational angle) $\Phi$ are described by
$\dot{\chi}^{A}$ $=$ $(G^{-1})^{A}_{B}\dot{I}^{B},$ (15)
@@
where a dot ($\cdot$) over variable denotes derivative with respect to $t$ and $G^{A}_{B}\equiv(\partial I^{A}/\partial\chi^{B})$ is the Jacobian for the transformation from $\chi^{A}$ to $I^{A}$.
@@ II.2 Spherical orbits
A spherical orbit is a bound orbit with $e=0$ (or equivalently, $r_{p}=r_{a}$), and hence is characterized by two orbital parameters, $\{p,\iota\}$. The parameters, $\{E,L,C\}$, are not independent for spherical orbits. In this work, we choose $\{E,L\}$ from them as a set of independent parameters.
For the later convenience, we introduce $x\equiv(M\omega_{\varphi})^{1/3}$ and $Y\equiv\cos\iota$ instead of $p$ and $\iota$, respectively. Since $x$ corresponds to the characteristic velocity in the motion, it is used as the parameter of the PN expansion. By using the set of $\{x,Y\}$, the equations of orbital evolution, Eqs. (15) and (16), for spherical orbits, are reexpressed as
@@
The specific energy and angular momentum for a spherical orbit at the leading order of the mass ratio are given in the form of the PN expansion  [^3] Although $E$ and $L$ (and also $G^{A}_{B}$) are known exactly, we treat their PN-expanded forms as in the standard PN calculation, adjusting to the previous works [37, 38]. as
@@
The superscript $(n)$ in the left-hand side of the above equations means that the formula includes the terms up to $n$-th order terms of $x$ relative to the Newtonian order  [^4] We consider the Newtonian order of the energy is $E^{(0)}=1-x^{2}/2$ although we do not employ this as the overall factor of the PN expansion. The term like $x^{n}(\ln x)^{k}$ is treated as $O(x^{n})$.. For example, the 2.5PN ($n=5$) energy and angular momentum are
$E^{(5)}(x,Y)$ $=$ $1-\frac{1}{2}x^{2}\left\{1-\frac{3}{4}x^{2}-\left(\frac{4}{3}-4Y\right)qx^{3}-\left[\frac{27}{8}-\left(\frac{1}{2}+Y-\frac{5}{2}Y^{2}\right)q^{2}\right]x^{4}+(2+6Y)qx^{5}\right\},$ (22)
$L^{(5)}(x,Y)$ $=$ $\frac{Y}{x}\left\{1+\frac{3}{2}x^{2}+\left(\frac{2}{3}-4Y\right)x^{3}+\left[\frac{27}{8}-\left(\frac{1}{4}+\frac{Y}{2}-\frac{7}{4}Y^{2}\right)q^{2}\right]x^{4}-(1+6Y)x^{5}\right\},$ (23)
@@
where $\mu$ denotes the mass of a point mass. In the same way of Eqs. (20) and (21), the superscript $(n)$ means the truncation order. It should be noted that the coefficients, $\tilde{\dot{E}}_{k}$ and $\tilde{\dot{L}}_{k}$, include some terms with the power of $\ln x$ for $k\geq 6$. We have obtained the coefficients for $k\leq 24$ (up to 12PN order) to date and confirmed that they are consistent with the $e\to 0$ (circular) limit of Ref. [20] and the $Y\to 1$ (equatorial) limit consistent with Ref. [22]. We have also confirmed that the 12PN formulas are consistent with the numerical results [14, 15] within numerical precision. In Appendix B, we show $\tilde{\dot{E}}_{k}$ and $\tilde{\dot{L}}_{k}$ for $k\leq 9$ explicitly (the higher order terms are publicly available online [31]).
@@ III Post-Newtonian templates of the phase evolution
The PN approximation provides various expressions of the phase evolution with different variables and functional forms while keeping the same PN order [36, 37, 38]. In this section, we present several PN families of the phase evolution of quasi-spherical orbits, TaylorT1, TaylorT2, TaylorT3, and TaylorT4 in the time domain  [^5] We do not consider the TaylorEt approximant because it is shown that the performance is not good compared to the other PN families for non-spinning, circular case [37, 38]. In the Kerr case, the specific energy depends on the odd power of $x$ and therefore the expansion of $x$ with respect to $\zeta=-2E$ used in TaylorEt includes the half-integer powers, which makes the approximant more complicated. This is another reason for not considering the TaylorEt in this paper., and TaylorF1, TaylorF2 in the frequency domain, in a similar manner to Refs. [36, 37, 38]. In addition, we present subfamilies of TaylorT1 and TaylorT4, which appear because the number of the orbital parameters increases compared to circular orbits.
@@ III.1 TaylorT1
Substituting the PN formulas given by Eqs. (20)–(25) to $E$, $L$, $\dot{E}$ and $\dot{L}$ in the right-hand sides of Eqs. (17) and (18), we obtain
@@
$\dot{\Phi}$ $=$ $\frac{x^{3}}{M},$ (28)
where $(G_{n})^{A}_{B}=\partial(E^{(n)},L^{(n)})/\partial(x,Y)$. Note that $(G_{n})^{A}_{B}$ is expanded in the PN form but the inverse, $(G_{n}^{-1})^{A}_{B}$, is not. Therefore, the right-hand sides of Eqs. (26) and (27) are expressed by rational functions of $x$, $\ln x$ and $Y$ although we do not show them explicitly because of their complexity.
By analogy of the circular orbit case, we refer to the numerical solution of Eqs. (26)–(28) as the TaylorT1 approximant and represent the solution with respect to $\Phi$ by $\Phi_{\textrm{T1}}^{(n)}$.
@@ III.2 TaylorT4
The difference between the TaylorT4 and TaylorT1 approximants  [^6] Historically, TaylorT1, TaylorT2, and TaylorT3 were developed first in Ref. [36], and then TaylorT4 was introduced to compare numerical relativity simulations for comparable-mass binary black holes in Ref. [39]. is whether we perform series expansion of the right-hand sides of Eqs. (26) and (27) with respect to $x$ or not. Expanding the right-hand sides of Eqs. (26) and (27) with respect to $x$, we obtain
@@
It should be noted that, since the PN formulas of the first and second terms in the right-hand side of Eq. (27) are given as
@@
the first two leading terms of them are canceling each other out. Hence, the leading term in the PN expansion of $\dot{Y}$ is considered as the 1.5PN contribution of the adiabatic evolution.
The TaylorT4 approximant is obtained by solving Eqs. (29)–(33) numerically. We represent the approximant with respect to $\Phi$ by $\Phi_{\textrm{T4}}^{(n)}$.